OpenStudy (anonymous):

Solve 152x = 36. Round to the nearest ten-thousandth. A. 0.6616 B. 2.6466 C. 1.7509 D. 1.9091

OpenStudy (anonymous):

divide both sides of the equation by 152....

OpenStudy (anonymous):

sorry it's suppose to be\[15^2x=36\]

OpenStudy (anonymous):

it's 15^2x=36

OpenStudy (luigi0210):

Darn you I have to start over

OpenStudy (anonymous):

so 2x is the exponent

OpenStudy (luigi0210):

Yea.. righyt?

OpenStudy (anonymous):

@rajee_sam @ajprincess

OpenStudy (anonymous):

yup, do u know how to do it?

OpenStudy (anonymous):

Take Log to both sides

OpenStudy (anonymous):

\[\Large 15^{2x}=36 \] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Take Log to Both sides ,it will be : \[\Large 15^{2x}=36 \] \[\Large (2x)Log15=Log36\] \[\Huge \color{MidnightBlue}{\bf 2x=\frac{Log36}{Log15}}\] I 'll let find the answer from here .

OpenStudy (anonymous):

it's A 0.6616